Write the equation of the line passing through each of the given pairs of points. Write your
result in slope-intercept form, where possible.
(-1, 3) and (4, -2)

Answer by checkley71(8403)

FIND THE SLOPE (Y2-Y1)/(X2-X1)
(-2-3)/(4+1)
-5/5=-1 FOR THE SLOPE. NOW REPLACE X & Y WITH ONE SET OF POINTS IN THE LINE EQUATION TO FIND THE Y INTERCEPT (b)
Y=mX+b
3=-1*-1+b
3=1+b
b=3-1
b=2
THUS THE EQUATION FOR THIS LINE IS
Y=-X+2